# Box Confidence Depth: simulation-based inference with hyper-rectangles

### v1

@@ 1 Introduction
This study focuses on frequentist inference, targeting the construction of calibrated confidence intervals and regions across simulation-based models and non-standard regularity conditions. The proposed approach provides a unified strategy for inference that seamlessly accommodates both univariate and multivariate parameters. This is achieved by means of a depth function (Liu, 1990), that allows defining nested confidence sets across all confidence levels, offering researchers a comprehensive visualization of parametric uncertainty. A significant aspect of the proposed methodology is its ability to operate without requiring data to be necessarily reduced to a scalar summary statistic. Raw data can be utilized directly, enhancing the flexibility and automation of the inference process, and inference from diverse test statistics, linked to model-specific information, can be combined in a natural manner. As a byproduct of the procedure, the method also yields consistent point estimators for model parameters.
The rest of the paper is organized as follows. Section 2 reviews recent developments in simulation-based inference. Section 3 outlines the sampling methodology used to build the Confidence Depth, discusses its theoretical underpinnings, and addresses some computational aspects. Section 4 discusses various examples from either classical models such as Generalized Linear Models (GLMs), as well as models from the field of Likelihood Free Inference (LFI) and reports simulation studies. A discussion is provided in Section 5.
@@ 2 Simulation based inference
The most popular method for SBI in Bayesian inference is Approximate Bayesian Computation (ABC), introduced by Rubin (1984) and further developed by Pritchard et al. (1999). ABC aims to generate datasets that mimic the observed sample using as proposals for $\theta$ draws from the prior distribution. Parameter values that generate synthetic observations closely matching the real observation, up to a certain tolerance $\varepsilon$, i.e. $d(y,y^{\text{obs}})<\varepsilon$, are retained. The distance $d(\cdot,\cdot)$ between pseudo and actual data is generally assessed on a set of summary statistics that are informative for the model. Intuitively, if the synthetic data match the observed data, the model parameters used in the simulations are plausible for the model under consideration and in turns, they are associated to higher likelihood function. Several enhancements to the basic ABC algorithm have been proposed over time, see Marjoram et al. (2003); Marin et al. (2012); Del Moral and Murray (2015); Frazier et al. (2018); Bernton et al. (2019); Rotiroti and Walker (2024) and references therein. The approximation in ABC is considered non parametric, as the shape of the likelihood and the posterior is not specified but obtained by rejection Monte Carlo. Parametric approximations of likelihood functions (and posteriors) in simulation-based settings have seen significant advancements in recent years, largely due to the growing influence of Machine Learning and Deep Learning techniques. Two prominent approaches are Bayesian Synthetic likelihood (Wood, 2010; Price et al., 2018; Frazier et al., 2023), which employs conditional density estimators based on a multivariate Gaussian model and the family of Neural Posterior Estimation methods (Rezende and Mohamed, 2015; Papamakarios, Sterratt and Murray, 2019) employing more flexible conditional density estimators, as normalizing flows which better suited for high-dimensional data and complex models. Machine Learning methods have been heavily employed in Neural Ratio Estimation (NRE) (Hermans, Begy and Louppe, 2020; Thomas et al., 2022) that estimates the ratio between the likelihood $p(y|\theta)$ and data marginal $p(y)$, that is $r(y,\theta)=p(y|\theta)/p(y)$, by training a classifier to distinguish datasets generated from the conditional and the marginal model.
@@ 3 Box-Confidence Depth
The proposed method consists in drawing $\theta^{*}$ from $\pi(\theta)$ and, for each $\theta^{*}$, generate two pseudo-samples from the model ${p(y|\theta^{*})}$, denoted as $y^{*1}$ and $y^{*2}$. Summary statistics $t^{*1}$ and $t^{*2}$ are then computed from $y^{*1}$ and $y^{*2}$, respectively, each of dimension $d$. The proposal $\theta^{*}$ is accepted if the observed summary statistic $t^{\text{obs}}$, computed from the actual observed data $y^{\text{obs}}$, falls within a region defined by $t^{*1}$ and $t^{*2}$. This region can be conceptualized as a $d-$dimensional hyper-rectangle, called Box and denoted as $\mathcal{B}_{t}^{*}$, in the space of summary statistics, with $t^{*1}=(t_{1}^{*1},\ldots,t_{d}^{*1})$ and $t^{*2}=(t_{1}^{*2},\ldots,t_{d}^{*2})$ defining its edges, i.e.
@@
The procedure described seeks to learn a data-dependent distribution over the parameter space non-parametrically, utilizing a rejection algorithm, as in ABC, instead of assuming a closed-form model for summary statistics. It differs from ABC because it focuses on establishing an ordering in the sample space to determine confidence level measure over the parameter space, which is consistent with the frequentist view on confidence and uncertainty. In particular, the acceptance rule does not make use of a distance function and a threshold parameter, but it is based on a series of inequalities. Note that, since establishing a meaningful ordering in the sample space becomes challenging in presence of multivariate data and summary statistics, the procedure in practice utilizes a measure of centrality of observed data with respect to simulated data, which corresponds to an ordering from the center outwards.
@@
$\mathcal{CD}^{\text{box}}(\theta):\Theta\mapsto R^{+},$
@@ 3.1 Scalar-scalar case
To formalize the properties of the function $\mathcal{CD}^{\text{box}}(\theta)$ in relation to confidence intervals and frequentist tests, it is useful to initially consider the scenario where $\theta\in\mathrm{R}$ and $d=1$. In this case, the Box reduces to an interval with endpoints $t(y^{*}_{1})$ and $t(y^{*}_{2})$ and the proposed $\theta^{*}$ is accepted if and only if $t(y^{\text{obs}})\in[t^{*(1)},t^{*(2)}]$.
@@
The statistic $t:\mathcal{Y}\mapsto\mathcal{T}\in\mathrm{R}$ is one-dimensional and $0<Var(t(y)|\theta)<\infty$ for $\pi$-almost all $\theta$.
The assumption that $Var(t(y)|\theta)>0$ for $\pi$-almost all $\theta$ ensures that the intervals of the form $[t^{(1)},t^{(2)}]$ have positive probability of being non-empty.
@@
where the function $F_{t}(t^{\text{obs}}|\theta)$ is the cumulative density function of $t(y)$, computed on the value of the observed summary statistic $t^{\text{obs}}$ and with $k$ of the same order of the machine tolerance.
@@
For a scalar parameter $\theta$, under Assumption 1 the Box-Confidence Depth is
@@
Statement $i)$ follows immediately from the definition of $F_{t}$ in Lemma 3.1. For $ii)$, in general, a function $H(y,\theta)$ on $\mathcal{Y}\times\Theta\rightarrow[0,1]$ is a CD for $\theta$ if (see e.g. Xie and Singh, 2013): a) for each given $y\in\mathcal{Y},H(\cdot)$ is a cumulative distribution function on $\Theta$; b) at $\theta=\theta_{0}$, $H(y^{\text{obs}},\theta_{0})$, as a function of the sample $y^{\text{obs}}$, follows a $\text{Uniform}[0,1]$ distribution. By construction $F_{t}(t^{\text{obs}}|\theta_{0})\in[0,1]$, furthermore, if $t$ is stochastically increasing in $\theta$, then $F_{t}(t^{\text{obs}}|\theta^{\prime})<F_{t}(t^{\text{obs}}|\theta^{\prime\prime})$ for $\theta^{\prime\prime}>\theta^{\prime}.$ By properties of the $p$-value function and for $u\in[0,1]$ $Pr(F_{t}(t^{\text{obs}}|\theta)<u)=Pr(t^{\text{obs}}<F_{t}^{-1}(u|\theta))$ is constant when $t^{\text{obs}}$ is drawn from $F_{t}(\cdot|\theta)$. ∎
@@
[…]
@@
Note that if the proposal $\pi(\theta)$ is centered on the confidence median, and the function $\mathcal{CD}^{\text{box}}(\theta)$ is symmetric, then the expected acceptance probability is 1/4. Note that if the proposal $\pi(\theta)$ is centered on the confidence median, and the function $\mathcal{CD}^{\text{box}}(\theta)$ is symmetric, then the expected acceptance probability is 1/4. This can be regarded as a practical guideline to tune the proposal.
@@
Input: Proposal distribution $\pi(\theta)$, number of iterations $R$, summary statistic function $t(\cdot)$, observed statistic $t^{\text{obs}}=t(y^{\text{obs}})$
Output: Accepted samples $\theta^{*}$
for $j\leftarrow 1$ to $R$ do
       Sample $\theta^{*}_{j}\sim\pi(\theta)$;
       Sample $y_{j}^{*1},y_{j}^{*2}\sim f(y;\theta_{j}^{*})$;
       Compute $t_{j}^{*1}=t(y_{j}^{*1})$,  $t_{j}^{*2}=t(y_{j}^{*2})$;
       if $t^{\text{obs}}\in\mathcal{B}(t_{j}^{*1},t_{j}^{*2})$ then
             Accept $\theta_{j}^{*}$;
       end if
end for
return Accepted samples $\theta^{*}$
Algorithm 1 Accept-Reject Box-CD
@@ 3.2 Relation to confidence intervals
[…]
@@
Consider the piecewise linear function $g(F_{t})=-|F_{t}-0.5|\in[-0.5,0]$. From its definition, the set $C_{1-\alpha}$ can be written as $\{\theta|$ $g(F_{t})\geq-0.5+\alpha/2\}$. Since the image of $g$ is $[-0.5,0]$ and the function is linear in $F_{t}$, if ${g(F_{t})}\geq-0.5+\alpha/2$ then ${g(F_{t})}\geq Q_{g}(\alpha)$. Applying the monotone transformation $h(F_{t})=2g(F_{t})\text{sign}(F_{t}-0.5)+F_{t}=2(0.5-F_{t})+F_{t}=F_{t}(1-F_{t})$ which is order preserving, it directly follows that $Q_{\mathcal{CD}^{\text{box}}(\theta)}(\alpha)={g(F_{t})}$. This concludes the proof as $C_{1-\alpha}$ can be written as $\{\theta|$ $\mathcal{CD}^{\text{box}}(\theta)\geq Q_{\mathcal{CD}^{\text{box}}(\theta)}(\alpha)\}$. ∎
@@ 3.3 Mutivariate case: center-outward ordering
$\text{HD}_{1}=2\min\{Pr(Y\leq y^{\text{obs}}\mid\theta),Pr(Y\geq y^{\text{obs}}\mid\theta)\}.$
@@
For two parameter points $\theta^{*}$ and $\theta^{**}$ within $\Theta$ and with their corresponding random Boxes $\mathcal{B}_{t}^{*}$, $\mathcal{B}_{t}^{**}$ , it holds that
@@ 3.4 Relation to Confidence sets and properties of the Box-CD function
We can generalize Lemma 3.4 to the general case of confidence sets assuming $\theta\in\mathrm{R}^{p}$. Define $M=\max_{\theta}\mathcal{CD}^{\text{box}}(\theta)$.
@@
$Pr_{\theta_{0}}(\mathcal{CD}^{\text{box}}(\theta)>{Q_{\mathcal{CD}}}(\alpha))=\alpha,$
@@
Define the transformation $w(t):t\in\mathrm{R^{d}}\mapsto w\in\mathrm{R^{d}}$ as a bijective, component-wise function, where $w_{j}(t):t_{j}\mapsto w_{j}$ represents the monotonic transformation applied specifically to the $j$-th component of the statistic $t$. Then $w_{j}^{\prime}>w_{j}\Leftrightarrow\{t_{j}^{\prime}>t_{j}\text{ or }t_{j}^{\prime}<t_{j}\}$ for any $t_{j}\in\mathrm{R}.$ In particualr, given $t_{j}$ such that $t_{j}^{\prime}<t_{j}<t_{j}^{\prime\prime}$, it follows that $w_{j}^{\prime}<w_{j}<w_{j}^{\prime\prime}$ or $w_{j}^{\prime}>w_{j}>w_{j}^{\prime\prime}$ and $Pr(t_{j}\in\mathcal{B}_{t}^{*})=Pr(w_{j}\in\mathcal{B}_{w}^{*})$. ∎

### v2

@@ 1 Introduction
This study focuses on frequentist inference, targeting the construction of calibrated confidence intervals and regions across simulation-based models and non-standard regularity conditions. The proposed approach provides a unified strategy for inference that seamlessly accommodates both univariate and multivariate parameters. This is achieved by means of a depth function (Liu, 1990), that allows defining nested confidence sets across all confidence levels, offering researchers a comprehensive visualization of parametric uncertainty. A significant aspect of the proposed methodology is its ability to operate without requiring data to be necessarily reduced to a scalar summary statistic, as it is typically done in the frequentist framework. Raw data can be utilized directly, enhancing the flexibility and automation of the inference process. Similarly, inference from diverse test statistics, linked to model-specific information, can be combined in a natural manner. As a byproduct of the procedure, the method also yields consistent point estimators for model parameters.
The rest of the paper is organized as follows. Section 2 reviews recent developments in simulation-based inference. Section 3 outlines the sampling methodology used to build the Confidence Depth, discusses its theoretical underpinnings, and addresses some computational aspects, with particular emphasis on challenges and remedies related to the curse of dimensionality problem. Section 4 discusses various examples from either classical models such as Generalized Linear Models (GLMs), as well as models from the field of Likelihood Free Inference (LFI) and reports simulation studies. A discussion is provided in Section 5.
@@ 2 Simulation based inference
The most popular method for SBI in Bayesian inference is Approximate Bayesian Computation (ABC), introduced by Rubin (1984) and further developed by Pritchard et al. (1999). ABC aims to generate datasets that mimic the observed sample using as proposals for $\theta$ draws from the prior distribution. Parameter values that generate synthetic observations closely matching the real observation, up to a certain tolerance $\varepsilon$, i.e. $d(y,y^{\text{obs}})<\varepsilon$, are retained. The distance or divergence $d(\cdot,\cdot)$ between pseudo and actual data is generally assessed on a set of summary statistics that are informative for the model. Intuitively, if the synthetic data match the observed data, the model parameters used in the simulations are plausible for the model under consideration and in turns they are associated to higher likelihood function. Several enhancements to the basic ABC algorithm have been proposed over time, see Marjoram et al. (2003); Marin et al. (2012); Del Moral and Murray (2015); Frazier et al. (2018); Bernton et al. (2019); Rotiroti and Walker (2024) and references therein. The approximation in ABC is considered non parametric, as the shape of the likelihood and the posterior is not specified but obtained by rejection Monte Carlo. Parametric approximations of likelihood functions (and posteriors) in simulation-based settings have seen significant advancements in recent years, largely due to the growing influence of Machine Learning and Deep Learning techniques. Two prominent approaches are Bayesian Synthetic likelihood (Wood, 2010; Price et al., 2018; Frazier et al., 2023), which employs conditional density estimators based on a multivariate Gaussian model and the family of Neural Posterior Estimation methods (Rezende and Mohamed, 2015; Papamakarios, Sterratt and Murray, 2019) employing more flexible conditional density estimators, as normalizing flows which better suited for high-dimensional data and complex models. Machine Learning methods have been heavily employed in Neural Ratio Estimation (NRE) (Hermans, Begy and Louppe, 2020; Thomas et al., 2022) that estimates the ratio between the likelihood $p(y|\theta)$ and data marginal $p(y)$, that is $r(y,\theta)=p(y|\theta)/p(y)$, by training a classifier to distinguish datasets generated from the conditional and the marginal model.
@@ 3 Box-Confidence Depth
The proposed method consists in drawing $\theta^{*}$ from $\pi(\theta)$ and, for each $\theta^{*}$, generate two pseudo-samples from the model $p(y|\theta^{*})$}, denoted as $y^{*1}$ and $y^{*2}$. Summary statistics $t^{*1}$ and $t^{*2}$, each of dimension $d$, are then computed from $y^{*1}$ and $y^{*2}$, respectively. The proposal $\theta^{*}$ is accepted if the observed summary statistic $t^{\text{obs}}$, computed from the actual observed data $y^{\text{obs}}$, falls within a region defined by $t^{*1}$ and $t^{*2}$. This region can be conceptualized as a $d$-dimensional hyper-rectangle, called Box and denoted as $\mathcal{B}_{t}^{*}$, in the space of summary statistics, with $t^{*1}=(t_{1}^{*1},\ldots,t_{d}^{*1})$ and $t^{*2}=(t_{1}^{*2},\ldots,t_{d}^{*2})$ defining its edges, i.e.
@@
Input: Proposal distribution $\pi(\theta)$, number of iterations $R$, summary statistic $t(\cdot)$, observed statistic $t^{\text{obs}}=t(y^{\text{obs}})$
Output: Accepted samples $\theta^{*}$
for $j\leftarrow 1$ to $R$ do
   Sample $\theta^{*}_{j}\sim\pi(\theta)$;
   Sample $y_{j}^{*1},y_{j}^{*2}\sim p(y|\theta_{j}^{*})$;
   Compute $t_{j}^{*1}=t(y_{j}^{*1})$,  $t_{j}^{*2}=t(y_{j}^{*2})$;
 if $t^{\text{obs}}\in\mathcal{B}_{t}^{*}$ then
      Accept $\theta_{j}^{*}$;
   end if
end for
return Accepted samples $\theta^{*}$
Algorithm 1 Accept-Reject Box-CD
@@
The procedure described seeks to learn a data-dependent distribution over the parameter space non-parametrically, utilizing a rejection algorithm, as in ABC, instead of assuming a closed-form model for summary statistics. However, it deeply differs from ABC as the inclusion criterion is not a distance function or a divergence. In particular, distances or divergences employed in ABC satisfy the identity of indiscernibles property, i.e. $d(a,b)=0\iff a=b$. Thinking about the proposed criterion illustrated in Algorithm 1 as a discrepancy, it still may be zero even if the simulated data don’t perfectly align with the observed sample. Conversely, at least ideally in ABC, as the tolerance or threshold parameter narrows, the pseudo-data must precisely match the observed data. The concept of matching simulations to observed data aligns with the idea of conditioning, which is a fundamental aspect of ABC. In contrast, the notion of ordering by using a series of inequalities in the sample space conforms to frequentist reasoning.
Note that, since establishing a meaningful ordering in the sample space becomes challenging in presence of multivariate data and summary statistics, the procedure in practice utilizes a measure of centrality of observed data with respect to simulated data, which corresponds to an ordering from the center outwards.
@@
$\mathcal{CD}^{\text{box}}(\theta):\Theta\mapsto{\color[rgb]{0,0,0}\mathbb{R}^{+}},$
@@
As a final remark, observe that the procedure can be generalized to consider more than two replicas of the data. This extension is discussed in Section 3.5, after detailing the method’s properties and the main results.
@@ 3.1 Scalar-scalar case
To formalize the properties of the function $\mathcal{CD}^{\text{box}}(\theta)$ in relation to confidence intervals and frequentist tests, it is useful to initially consider the scenario where $\theta\in{\color[rgb]{0,0,0}\mathbb{R}}$ and $d=1$. In this case, the Box reduces to an interval with endpoints $t^{*1}=t(y^{*}_{1})$ and $t^{*2}=t(y^{*}_{2})$ and the proposed $\theta^{*}$ is accepted if and only if $t(y^{\text{obs}})\in[t^{*(1)},t^{*(2)}]$.
@@
The statistic $t:\mathcal{Y}\mapsto\mathcal{T}\in\mathbb{R}$ is one-dimensional and $0<Var(t(y)|\theta)<\infty$ for $\pi$-almost all $\theta$.
The assumption that $Var(t(y)|\theta)>0$ for $\pi$-almost all $\theta$ ensures that the intervals of the form $t^{(1)},t^{(2)}]$ have positive probability of being non-empty.
@@
where $F_{t}(t^{\text{obs}}|\theta)$ is the cumulative distribution function of $t(y)$, computed at the value of the observed summary statistic $t^{\text{obs}}$ and with $k$ of the same order of the machine tolerance.
@@
For a scalar parameter $\theta$, under Assumption 1, the Box-Confidence Depth is
@@
Statement $i)$ follows immediately from the definition of $F_{t}$ in Lemma 3.1. For ii)ii), in general, a function $H(y,\theta)$ on $\mathcal{Y}\times\Theta\rightarrow[0,1]$ is a CD for $\theta$ if (see e.g. Xie and Singh, 2013): a) for each given $y\in\mathcal{Y},H(\cdot)$ is a cumulative distribution function on $\Theta$; b) at $\theta=\theta_{0}$, $H(y^{\text{obs}},\theta_{0})$, as a function of the sample $y^{\text{obs}}$, follows a $\text{Uniform}[0,1]$ distribution. By construction $F_{t}(t^{\text{obs}}|\theta_{0})\in[0,1]$, furthermore, if $t$ is stochastically increasing in $\theta$, then $F_{t}(t^{\text{obs}}|\theta^{\prime})<F_{t}(t^{\text{obs}}|\theta^{\prime\prime})$ for $\theta^{\prime\prime}>\theta^{\prime}.$ By properties of the $p$-value function and for $u\in[0,1]$ $Pr(F_{t}(t^{\text{obs}}|\theta)<u)=Pr(t^{\text{obs}}<F_{t}^{-1}(u|\theta))$ is constant when $t^{\text{obs}}$ is drawn from $F_{t}(\cdot|\theta)$. ∎
@@
By definition $\hat{\theta}=\underset{\theta}{\arg\max}\>F_{t}(t^{\text{obs}}|\theta)[1-F_{t}(t^{\text{obs}}|\theta)].$ Since $F_{t}(t^{\text{obs}}|\theta)\in[0,1]$, the expression is maximum when the function x(1x)x(1-x) is maximum with $x$ in $[0,1]$, which is $F_{t}(t^{\text{obs}}|\hat{\theta})=0.5$. Then, applying $F_{t}$ to both sides of the inequality of Equation (1) it follows that $Pr_{\theta_{0}}(F_{t}(t^{\text{obs}}|\hat{\theta})<F_{t}(t^{\text{obs}}|\theta_{0}))=$ $Pr_{\theta_{0}}(0.5<F_{t}(t^{\text{obs}}|\theta_{0}))=1/2.$
Remark 1.
@@
Note that if the proposal $\pi(\theta)$ is centered on the confidence median, and the function $\mathcal{CD}^{\text{box}}(\theta)$ is symmetric, then the expected acceptance probability is 1/4. This can be regarded as a practical guideline to tune the proposal.
@@ 3.2 Relation to confidence intervals
Let us write $F_{t}(\theta)$ shortly for $F_{t}(t^{\text{obs}}|\theta)$. Define an equi-tailed confidence interval of size $1-\alpha$ as $C_{1-\alpha}=\{\theta|F_{t}(\theta)>\alpha/2\text{\;and\>}F_{t}(\theta)<(1-\alpha/2)\}$. Denote by $Q_{Z}(\cdot)$ the function s.t. for $p\in(0,1)$ and $Z,\zeta\in(0,1)$, $Q_{Z}(p)=\zeta$ if $\int 1_{Z<\zeta}dZ=p$.
@@
Consider the piecewise linear function $g(F_{t})=-|F_{t}-0.5|\in[-0.5,0]$. From its definition, the set $C_{1-\alpha}$ can be written as $\{\theta|$ $g(F_{t})\geq-0.5+\alpha/2\}$. Since the image of $g$ is [0.5,0][-0.5,0] and the function is linear in $F_{t}$, if ${g(F_{t})}\geq-0.5+\alpha/2$ then ${g(F_{t})}\geq Q_{g}(\alpha)$. Applying the monotone transformation $h(F_{t})=2g(F_{t})\text{sign}(F_{t}-0.5)+F_{t}=2(0.5-F_{t})+F_{t}=F_{t}(1-F_{t})$ which is order preserving, it directly follows that $Q_{\mathcal{CD}^{\text{box}}(\theta)}(\alpha)={g(F_{t})}$. This concludes the proof as $C_{1-\alpha}$ can be written as $\{\theta|$ $\mathcal{CD}^{\text{box}}(\theta)\geq Q_{\mathcal{CD}^{\text{box}}(\theta)}(\alpha)\}$. ∎
@@ 3.3 Mutivariate case: center-outward ordering
$\text{HD}_{1}=2\min\{Pr_{\theta}(Y\leq y^{\text{obs}}),Pr(Y\geq y^{\text{obs}})\}.$
@@
For two parameter points $\theta^{*}$ and $\theta^{**}$ within $\Theta$ and with their corresponding random Boxes $\mathcal{B}_{t}^{*}$, $\mathcal{B}_{t}^{**}$, it holds
@@ 3.4 Relation to Confidence sets and properties of the Box-CD function
We can generalize Lemma 3.4 to the general case of confidence sets assuming $\theta\in\mathbb{R}^{p}$. Define $M=\max_{\theta}\mathcal{CD}^{\text{box}}(\theta)$.
@@
$Pr_{\theta_{0}}(\mathcal{CD}^{\text{box}}(\theta)>{Q_{\mathcal{CD}^{\text{box}}}}(\alpha))=\alpha,$
@@
Define the transformation $w(t):t\in\mathbb{R^{d}}\mapsto w\in\mathbb{R^{d}}$ as a bijective, component-wise function, where $w_{j}(t):t_{j}\mapsto w_{j}$ represents the monotonic transformation applied specifically to the $j$-th component of the statistic $t$. Then $w_{j}^{\prime}>w_{j}\Leftrightarrow\{t_{j}^{\prime}>t_{j}\text{ or }t_{j}^{\prime}<t_{j}\}$ for any $t_{j}\in\mathbb{R}.$ In particualr, given $t_{j}$ such that $t_{j}^{\prime}<t_{j}<t_{j}^{\prime\prime}$, it follows that $w_{j}^{\prime}<w_{j}<w_{j}^{\prime\prime}$ or $w_{j}^{\prime}>w_{j}>w_{j}^{\prime\prime}$ and $Pr(t_{j}\in\mathcal{B}_{t}^{*}|\theta)=Pr(w_{j}\in\mathcal{B}_{w}^{*}|\theta)$. ∎
3.5 Efficiency and optimality
In the Box-CD framework, coverage validity and type-I error control are guaranteed, while the width of the confidence sets reflects the amount of information preserved in the summary statistics. For instance, if a sample of size $n$ is processed by computing statistics on only a fraction of the available observations, the effective information content decreases, which in turn leads to wider confidence intervals.
Under classical regularity conditions—namely independent and identically distributed data from a regular parametric family, smoothness of the likelihood, existence of sufficient statistics (often complete in exponential families), finite Fisher information, and the Monotone Likelihood Ratio (MLR) property—confidence distributions for a scalar parameter achieve optimality. In particular, they yield confidence intervals with the shortest possible expected length while maintaining the prescribed coverage probability, as their construction aligns with the theory of uniformly most powerful tests. Within this framework, Box-CD–based intervals correspond to equi-tailed intervals derived from the distribution of the pivotal quantity and the confidence distribution is simply the parameter-space representation of the pivot’s distribution, and Box-CD–based intervals emerge as the equi-tailed confidence intervals obtained from this construction. When constructed from multiple summary statistics, the Box-CD can be directly compared to likelihood ratio methods in terms of efficiency.
Theorem 3.8.
CD-based test versus Likelihood Ratio Test (LRT) Let $y=(y_{1},\dots,y_{n})$ be i.i.d. from a family of distributions $\{p(y\mid\theta):\theta\in\Theta\}$. Suppose that for each $i$, the family $\{p(y_{i}\mid\theta):\theta\in\Theta\}$ has the MLR property for a given statistic $t(y_{i})$. Define the likelihood ratio for testing $H_{0}:\theta=\theta_{0}$ versus $H_{1}:\theta\neq\theta_{0}$ as
$\lambda(y)=L(\theta_{0}\mid y)/L(\hat{\theta}\mid y),$
where $L(\theta\mid y)=\prod_{i=1}^{n}p(y_{i}\mid\theta)$ and $\hat{\theta}$ is the MLE under the full parameter space $\Theta$. The test rejects for small values of $\lambda(y)$ and we denote the corresponding critical values at significance level $\alpha$ by $c_{\alpha}$. Then:
  1. (i)
    For a single observation $y_{i}$, the test based on the Box-CD $\mathcal{CD}_{i}^{\mathrm{box}}(\theta)$, which rejects $H_{0}$ when
    $\mathcal{CD}_{i}^{\mathrm{box}}(\theta_{0})<\alpha,$
    is equivalent to the likelihood ratio test (LRT); that is, it has the same rejection region as the LRT applied to $y_{i}$.
  2. (ii)
    For the full sample $y=(y_{1},\dots,y_{n})$, the critical region of the LRT at level $\alpha$ is determined by $c^{*}_{\alpha}$, which is the smallest value among the coordinate-wise critical values $c_{i\alpha}$ for individual LRTs:
    $c^{*}_{\alpha}\;=\;\min_{1\leq i\leq n}c_{i\alpha}.$
  3. (iii)
    For any $\theta$ and $\alpha$, there exists an integer $k$ such that when $\mathcal{CD}^{\text{box}}(\theta_{0})=\alpha$ $\mathcal{CD}^{\text{box}}_{k}(\theta_{0})=\alpha_{k}$, with $\alpha_{k}\leq\alpha$.
As a consequence, the rejection region of the CD-based test contains that of the LRT at the same nominal significance level $\alpha$. Therefore, the power function of the CD-based test satisfies
$\beta_{\mathrm{CD}}(\theta)\;\geq\;\beta_{\mathrm{LRT}}(\theta),\quad\text{for all }\theta\in\Theta_{1},$ (2)
where $\Theta_{1}$ denotes the parameter values under the alternative hypothesis. This inequality may be strict for some $\theta$, indicating that the CD-based test can achieve strictly higher power than the LRT in these cases. In other words, under the stated MLR and aggregation conditions, the CD-based test is uniformly at least as powerful as the LRT and may outperform it for certain alternatives.
Proof.
By the MLR property, each marginal Box-CD test $\mathcal{CD}_{i}^{\text{box}}(\theta)$ based on $y_{i}$ is equivalent to the LRT for that observation. In particular, for a single observation,
$\mathcal{CD}_{i}^{\text{box}}(\theta_{0})<\alpha_{i}\quad\Leftrightarrow\quad\lambda(y_{i})<c_{i\alpha},$
so their rejection regions coincide. For the full sample $y=(y_{1},\dots,y_{n})$, let us express the aggregated Box-CD in terms of conditional probabilities. For any two marginals $i$ and $j$:
$\mathcal{CD}^{\text{box}}(\theta_{0})=\frac{\Pr(t_{i}\in\mathcal{B}_{i}\text{ and }t_{j}\in\mathcal{B}_{j})}{M}=\frac{\Pr(t_{i}\in\mathcal{B}_{i})\,\Pr(t_{j}\in\mathcal{B}_{j}\mid t_{i}\in\mathcal{B}_{i})}{M},$
where $\mathcal{B}_{i}\text{ and }\mathcal{B}_{j}$ are marginal boxes and $M$ is the maximum of the unnormalized Box-CD as defined in Section 3.4. Since $\mathcal{CD}_{i}^{\text{box}}(\theta_{0})=\Pr(t_{i}\in\mathcal{B}_{i})/M_{i}<\alpha_{i}$ for some $i$, and because conditional probabilities satisfy $0\leq\Pr(t_{j}\in\mathcal{B}_{j}\mid t_{i}\in\mathcal{B}_{i})\leq 1$, and $M_{i}>M$, the aggregated Box-CD satisfies
$\mathcal{CD}^{\text{box}}(\theta_{0})\leq\mathcal{CD}_{i}^{\text{box}}(\theta_{0})<\alpha_{i}.$
Thus, the aggregated rejection region contains all marginal rejection regions, including the LRT rejection region. Regarding power, we observe that:
  • if marginal statistics are weakly correlated, i.e. $\Pr(t_{j}\in\mathcal{B}_{j}\mid t_{i}\in\mathcal{B}_{i})\approx\Pr(t_{j}\in\mathcal{B}_{j})$, the aggregated Box-CD behaves similarly to the LRT, and power is approximately equal;
  • if there is moderate conditional dependence, i.e. $\Pr(t_{j}\in\mathcal{B}_{j}\mid t_{i}\in\mathcal{B}_{i})>\Pr(t_{j}\in\mathcal{B}_{j})$, aggregation increases evidence against $H_{0}$, so the Box-CD test can achieve strictly higher power than the LRT.
Hence, for all $\theta\neq\theta_{0}$, Equation 2 holds, with strict inequality under moderate conditional dependence. ∎